# OneVL: One-Step Latent Reasoning and Planning with Vision-Language Explanation

### 04 May 2026 v2

@@ 6 Conclusion
At inference time, the decoders are discarded and the prefilled latent tokens enable single-pass trajectory generation at a latency matching answer-only prediction. A three-stage training pipeline is essential for realizing this compression: Stage 0 establishes meaningful latent representations, Stage 1 aligns the decoders against a stable latent space, and Stage 2 tightens the bottleneck from both sides through bidirectional joint optimization. Ablations show that skipping this curriculum causes the compression to collapse into memorization, with a catastrophic performance drop.
Empirically, OneVL achieves state-of-the-art results on NAVSIM, ROADWork, Impromptu, and APR1 while matching answer-only inference speed. Most notably, OneVL is the only latent CoT method to outperform explicit autoregressive CoT, providing direct evidence that tighter compression, when grounded in both linguistic and world-model supervision, yields more effective reasoning than verbose token-by-token generation. Finally, by appending a lightweight MLP regression head, OneVL retains competitive performance at only 5% of the original latency, pointing toward efficient real-world deployment.
@@
As shown in Section˜5.6, appending an MLP head can improve inference speed at some cost to performance, so the overall latency remains constrained by AR decoding. Bridging this gap, for instance, through parallel or non-autoregressive trajectory decoding, is a key step toward true real-time deployment in safety-critical driving systems. Second, extending the world model decoder to multi-camera inputs would enable 360-degree future-scene prediction and more comprehensive causal scene understanding, further strengthening the compression targets available to the visual latents. Third, the dual-modal explanation framework could enable novel human-machine interface designs in which drivers receive real-time visual and verbal justifications for the vehicle’s planning decisions. Beyond enhancing transparency and trust, this also has the potential to supply richer training signals for reinforcement learning, closing the loop between world model prediction and policy improvement.
@@ Acknowledgments
[…]
8 Appendix
8.1 Data Format Example
Each NAVSIM training sample in OneVL follows the structure below. The assistant response contains visual latent tokens (4), language latent tokens (2), and the trajectory answer:
User: <image> Front-view image of the driving scene. Command: MOVE FORWARD. Velocity: [1.5, 0.0]. Acceleration: [0.3, 0.0]. Historical trajectory: (0.0, 0.0, 0.0), (-0.75, 0.0, 0.0), (-1.5, 0.0, 0.0). [Task instruction to output reasoning and predicted trajectory…] Assistant: <|start-latent-vis|> <|latent-vis|> <|latent-vis|> <|latent-vis|> <|latent-vis|>
<|end-latent-vis|> <|start-latent|> <|latent|> <|latent|> <|end-latent|>
<answer>[0.75, 0.0, 0.0], [1.5, 0.0, 0.0], …</answer>
think_steps (label for $\mathcal{D}_{\text{l}}$): “The road ahead is clear. There are no vehicles or pedestrians that would require braking. I should maintain current speed and continue straight along the lane.” future_image_tokens (label for $\mathcal{D}_{\text{v}}$):
<|image_start|>…<|visual_token_XXXXXX|>…<|image_end|>
8.2 Training Configuration
We begin with self-supervised pretraining of the visual auxiliary decoder: the model learns to predict the next frame from ViT features alone, without the autoregressive backbone in the loop. This phase uses 13 04013\,040 optimizer steps and a global batch size of $256$, rather than a fixed epoch count. We then move to Stage 0, which warms up the main VLM while introducing latent tokens, followed by Stage 1, where the auxiliary decoders train with the main model frozen and the visual decoder becomes action-conditioned through visual latent tokens. Finally, Stage 2 performs joint end-to-end fine-tuning with all parts of the stack updated together; details are shown in Table 9.
8.3 CoT Annotation Construction
The chain-of-thought reasoning annotations for the ROADWork dataset are constructed using an in-house annotation pipeline developed specifically for work-zone driving scenarios. Unlike the NAVSIM annotations (sourced from AdaThinkDrive [78]), ROADWork CoT reasoning requires domain-specific annotation of:
  • Work-zone hazard identification: Detection and description of cones, barriers, temporary signage, and worker presence.
  • Non-standard lane interpretation: Reasoning about temporary lane markings, reduced lane widths, and merge configurations.
  • Speed and clearance reasoning: Justification of appropriate speed reduction and lateral clearance decisions specific to work zones.
The CoT reasoning annotations for the Impromptu VLA Dataset are constructed using a VLM-centric annotation pipeline with Chain-of-Thought prompting developed specifically for unstructured driving scenarios (the core corner cases for autonomous driving). Unlike annotations for conventional structured driving datasets, Impromptu VLA CoT reasoning requires domain-specific annotation of:
  • Unstructured scenario classification: Identification and categorization of four core challenging unstructured scenarios, including roads with unclear boundaries, temporary traffic rule changes, unconventional dynamic obstacles, and challenging road conditions.
  • Complex scene element perception: Detection and description of ambiguous road boundaries, temporary traffic facilities, non-standard dynamic obstacles, and adverse road/environmental conditions.
  • Non-standard driving behavior reasoning: Interpretation of temporary traffic rules, obstacle avoidance strategies, and justification of speed adjustment and lateral clearance decisions in unstructured environments.
  • Planning-oriented decision reasoning: Generation of ego-vehicle meta-action plans, end-to-end trajectory predictions, and textual rationales for driving maneuvers.
Since the CoT reasoning labels for APR1 are not released, we replicate these annotations ourselves. Specifically, we use the publicly released model checkpoint [^1] https://huggingface.co/nvidia/APR1-10B to reproduce the CoT labels for all training examples. Regarding the waypoint prediction, we applied a heuristic subsampling strategy to reduce the sequence from 64 to 8 points while ensuring the last points are retained, as 64 waypoints are too dense for autoregressive modeling (the original APR1 model employs flow matching). Furthermore, as the original paper does not release the official test set, we subsample 700 examples from the available video clips to construct our test set.
The following are four anonymized CoT examples:
NAVSIM Example
The right side of the lane where ego vehicle is located is close to the undrivable area, so I need to drive slightly to the left. I should pay more attention to a vehicle, located 18.42 meters ahead of ego vehicle and 0.49 meters to the left. It is driving forward in the same direction, and its motion state is moving fastly. Based on the understanding of the driving scene and the navigation information, the ego should maintain speed and keep lane.
ROADWork Example
critical_factors: An orange work-zone warning sign placed in the roadway ahead indicates construction activity and prompts caution and adjustment in path. A line of cones along the left side of the roadway creates a temporary channelization/taper that narrows the usable lane and guides vehicles away from the left edge. Parked vehicles tightly line the right curb, reducing lateral clearance and constraining the ego vehicle available path through the work area. Scaffolding/sidewalk construction along the right side suggests active work frontage and encourages keeping distance from the curbside area. cause:Taper or channelization speed adjustment and channelized lane keeping because cones taper the left side and parked vehicles along the right curb narrow the usable lane. decision:longitudinal:Taper or channelization speed adjustment lateral:Channelized lane keeping
Impromptu Example
<decision>longitudinal decision: stop, lateral decision: straight</decision><cause> The decision to stop the car is likely due to the presence of construction barriers and a pedestrian crossing the road, indicating a need to ensure safety and comply with traffic regulations.</cause>
APR1 Example
Turn right at the intersection since the right-turn traffic light is green.
8.4 LLM-as-Judge Evaluation Prompt
We employ a state-of-the-art proprietary VLM, gemini-3.1-flash-lite-preview [^2] https://ai.google.dev/gemini-api/docs/models/gemini-3.1-flash-live-preview as an automated evaluator for text CoT quality. For each sample, the judge receives the front-camera image alongside the following prompt (with the ground-truth and predicted CoT filled in):
You are an expert evaluator for Autonomous Driving Systems.
Your task is to evaluate the quality of a predicted driving Chain-of-Thought (CoT) against the Ground Truth CoT. You are also provided with the front-camera image for driving context.
Ground Truth CoT: “{gt_cot}”
Predicted CoT: “{pred_cot}”
Evaluation Criteria:
1. Perception (Distance & Location): Minor numerical differences (e.g., 29.3m vs 27.5m) are acceptable and should only incur slight penalties.
2. Motion State (Prediction): Identifying the correct state (e.g., ‘moving fastly’, ‘moving slowly’, ‘keep static’) is highly critical. A mismatch here is a severe safety hazard and must heavily reduce the score.
3. Ego Decision (Planning): The final action (e.g., ‘accelerate’, ‘decelerate’, ‘maintain speed’, ‘keep lane’) MUST match the Ground Truth. Any deviation is a critical safety error and should result in a massive penalty.
4. Language Fluency: Penalize minor grammatical errors.
Score the Predicted CoT on a scale of 0 to 100 based on the above criteria.
Output STRICTLY in the following JSON format without any markdown blocks, backticks, or extra text:
{"reasoning": "Briefly explain the penalties applied based on the criteria.", "score": <integer_between_0_and_100>}
The judge’s raw score (0–100) is normalized to [0, 1][0,\,1] by dividing by 100. We set the temperature to 0.1 to ensure scoring stability.
Hyperparameter Pre-training Stage 0 Stage 1 Stage 2
Steps 13 04013\,040
Epochs 2 1 5
Batch (global) 256256 64 64 64
Learning rate $1{\times}10^{-4}$ $4{\times}10^{-5}$ ${\times}10^{-4} 1×1$${\times}10^{-4}$
LR schedule Cosine Cosine Cosine Cosine
Optimizer AdamW AdamW AdamW AdamW
Precision BF16 BF16 BF16 BF16
Parallelism Zero-2 Zero-2 Zero-2 Zero-2
Trainable Vis aux decoder ViT, LLM, aligner Lang. & vis aux dec All
Frozen Main VLM
[…]
Table 9: Training hyperparameters for OneVL.
Traj. Pred. L2 Error (m)
$\circ$ Impromptu VLA [17]* 0.90 2.80 3.75 5.89 3.16
$\circ$ Impromptu VLA [17] 0.14 0.60 1.45 2.67 1.22
[…]
Table 10: Performance comparisons on the Impromptu benchmark [17]. We report trajectory prediction L2 error following the benchmark setting (meters; lower is better).
8.5 Reproducing Impromptu
To reproduce the optimal performance of the baseline model, we follow the official settings to conduct evaluations using both the training script released by Impromptu and the pre-trained checkpoint uploaded to Hugging Face, with the corresponding test results presented in Table 10. Concretely, the first row marked with an asterisk (*) reports the results evaluated on the author-provided checkpoint [^3] https://huggingface.co/aaaaaap/ImpromptuVLAModel/tree/main/7B_AD, and the second row shows the outcomes of the model trained from scratch via the official training script. As neither of these results matches the performance claimed in the original paper, we adopt the better-performing one as the representative result of Impromptu for comparative analysis in the main experiments.
8.6 NAVSIM qualitative examples
Figure 9 to 17 show additional NAVSIM examples compare a baseline answer-only prediction against OneVL, along with two future frames decoded from the visual auxiliary decoder, the text CoT from the language decoder. The key reasoning that affects the decision is highlighted in bold. Each plot overlays ground-truth (green) and predicted (red) trajectories on the front camera view.
8.7 Roadwork qualitative examples
Figure 18 to 19 show additional ROADWork examples compare a baseline answer-only prediction against OneVL, along with two future frames decoded from the visual auxiliary decoder, the text CoT from the language decoder. Each plot overlays ground-truth (green) and predicted (red) trajectories on the front camera view.
8.8 Impromptu qualitative examples
Figures 20 and 21 present the cases of the answer-only baseline and the OneVL model on the impromptu task. Green lines in the figures denote the ground-truth trajectories provided by the dataset, and red lines represent the trajectories predicted by different models. As observed in Figure 20, the CoT prompt guides the model to avoid large vehicles merging into the traffic flow, enabling OneVL to generate trajectories that stay away from large vehicles and are closer to the ground truth. Furthermore, Figure 21 shows that when passing through a Y-shaped intersection, the right-turn CoT prompt allows OneVL to output a reasonable right-turn trajectory that remains more centered in the lane. Comprehensive analysis demonstrates that the information provided by CoT prompts can help produce trajectories that are closer to the ground truth and better satisfy driving requirements, especially in scenarios that require changes in current driving behavior and trajectories, such as crossing intersections and interacting with surrounding vehicles.
Figure 9: NAVSIM qualitative example 1.
Figure 10: NAVSIM qualitative example 2.
Figure 11: NAVSIM qualitative example 3.
Figure 12: NAVSIM qualitative example 4.
Figure 13: NAVSIM qualitative example 5.
Figure 14: NAVSIM qualitative example 6.
Figure 15: NAVSIM qualitative example 7.
Figure 16: NAVSIM qualitative example 8.
Figure 17: NAVSIM qualitative example 9.
Figure 18: ROADWork qualitative example 1.
Figure 19: ROADWork qualitative example 2.
Figure 20: Impromptu qualitative example 1.
Figure 21: Impromptu qualitative example 2.
8.9 Alpamayo-R1 qualitative examples
Figure 22 presents the cases of the answer-only baseline and the OneVL model on the Alpamayo-R1 dataset.
Figure 22: Alpamayo-R1 qualitative example.

### 08 May 2026 v3

@@ 6 Conclusion
At inference time, the decoders are discarded, and the prefilled latent tokens enable single-pass trajectory generation at a latency matching answer-only prediction. A three-stage training pipeline is essential for realizing this compression: Stage 0 establishes meaningful latent representations, Stage 1 aligns the decoders against a stable latent space, and Stage 2 tightens the bottleneck from both sides through bidirectional joint optimization. Ablations show that skipping this curriculum causes the compression to collapse into memorization, with a catastrophic performance drop.
Empirically, OneVL achieves state-of-the-art results on NAVSIM, ROADWork, Impromptu, and Alpamayo-R1 while matching answer-only inference speed. Most notably, OneVL is the only latent CoT method to outperform explicit autoregressive CoT, providing direct evidence that tighter compression, when grounded in both linguistic and world-model supervision, yields more effective reasoning than verbose token-by-token generation. Finally, by appending a lightweight MLP regression head, OneVL retains competitive performance at only 5% of the original latency, pointing toward efficient real-world deployment.
@@
As shown in Section˜5.6, appending an MLP head can improve inference speed at some cost to performance, so the overall latency remains constrained by AR decoding.
Bridging this gap, for instance, through parallel or non-autoregressive trajectory decoding, is a key step toward true real-time deployment in safety-critical driving systems. Second, extending the world model decoder to multi-camera inputs would enable 360-degree future-scene prediction and more comprehensive causal scene understanding, further strengthening the compression targets available to the visual latents. Third, the dual-modal explanation framework could enable novel human-machine interface designs in which drivers receive real-time visual and verbal justifications for the vehicle’s planning decisions. Beyond enhancing transparency and trust, this also has the potential to supply richer training signals for reinforcement learning, closing the loop between world model prediction and policy improvement.
7 Contributions and Acknowledgments
Core Contributors
  • Jinghui Lu
  • Jiayi Guan
  • Zhijian Huang
  • Jinlong Li
  • Guang Li
  • Lingdong Kong
  • Yingyan Li
  • Han Wang
  • Shaoqing Xu
  • Yuechen Luo
  • Fang Li
  • Chenxu Dang
  • Junli Wang
  • Tao Xu
  • Jing Wu
  • Jianhua Wu
  • Xiaoshuai Hao
  • Wen Zhang
  • Tianyi Jiang
  • Kuiyuan Yang
  • Hangjun Ye
  • Long Chen
Contributors
  • Lingfeng Zhang
  • Lei Zhou
  • Yingbo Tang
  • Jie Wang
  • Yinfeng Gao
  • Xizhou Bu
  • Haochen Tian
  • Yihang Qiu
  • Feiyang Jia
  • Lin Liu
  • Yigu Ge
  • Hanbing Li
  • Yuannan Shen
  • Jianwei Cui
  • Hongwei Xie
  • Bing Wang
  • Haiyang Sun
  • Jingwei Zhao
  • Jiahui Huang
  • Pei Liu
  • Zeyu Zhu
  • Yuncheng Jiang
  • Zibin Guo
  • Chuhong Gong
  • Hanchao Leng
  • Kun Ma
  • Naiyan Wang
  • Guang Chen
Corresponding author.
@@ Acknowledgments
In this work, we utilize the Qwen3-VL-4B-Instruct [5] model and the following datasets: NAVSIM [20], ROADWork [34], Impromptu [17], and Alpamayo-R1 [105]. The authors confirm that the use of these resources is strictly for academic research purposes and have not been involved in any commercial activities.